By knocking the glass vase, a student chipped off 0.070 g square of glass. How many silicon dioxide molecules were chipped off the vase?
Answer #1

Moles of SiO2 = 0.070/60

Number of molecules = 0.070×6.022×1023/60

= 7.02×1020 molecules